One grandfather has three grandchildren, two of their age difference is 3, eldest child age is 3 times youngest child’s age and eldest child’s age is two times of sum of other two children. What is the age of eldest child?
Found 2 solutions by ankor@dixie-net.com, aditi.mca:
Answer by ankor@dixie-net.com(22740)

One grandfather has three grandchildren, two of their age difference is 3,
eldest child age is 3 times youngest child’s age and
eldest child’s age is two times of sum of other two children.
What is the age of eldest child?
:
Three grand-kids, x,y,z; z is the eldest
:
"eldest child age is 3 times youngest child’s age"
z = 3x
:
"and eldest child’s age is two times of sum of other two children.:
z = 2(x+y)
z = 2x + 2y
Replace z with 3x
3x = 2x + 2y
3x - 2x = 2y
x = 2y; but that make the y the youngest??? Don't think this can be done!!

Answer by aditi.mca(1)

z=3x
z=2(x+y)
z=2x+2y
3x=2x+2y.
therefore x=2y
and z-y=3
substitute z=3x
we get y=3,x=6 and z=18...z=18 is the answer